# Value-Free Policy Optimization via Reward Partitioning

### v2

@@ III Method
$\footnotesize\pi^{*}(y|x)=\frac{\pi_{\mathrm{ref}}(y|x)\exp\left(\frac{1}{\tau}r(x,y)\right)}{Z(x)},$ (15)
where the partition function $Z(x)$ is
$\footnotesize Z(x)=\sum_{y^{\prime}}\pi_{\mathrm{ref}}(y^{\prime}|x)\exp\left(\frac{1}{\tau}r(x,y^{\prime})\right).$ (16)
Current methods such as DRO learn both a policy $\pi_{\theta}$ and a value function $V_{\phi}$ parameterized by $\phi$, where the soft value function relates to the partition function as $V^{*}(x)=\tau\log Z(x)$. However, this joint learning introduces several difficulties: the value function $V_{\phi}$ can be unstable to approximate, the parameters $(\theta,\phi)$ are tightly coupled, and optimization becomes more complex.
@@
$\footnotesize r(x,y)-V^{*}(x)=\tau\log\frac{\pi^{*}(y|x)}{\pi_{\mathrm{ref}}(y|x)},$ (17)
@@
$\footnotesize\mathcal{L}_{\mathrm{RPO}}(\theta)=\frac{1}{2n}\sum_{i=1}^{n}\left(\log\frac{\pi_{\theta}(y_{i}|x_{i})}{\pi_{\mathrm{ref}}(y_{i}|x_{i})}-\frac{1}{\tau}\left(r_{i}-\widehat{V}(x_{i})\right)\right)^{2}.$ (21)
@@
$\footnotesize\nabla_{z_{\theta}(y|x)}\mathcal{L}_{\mathrm{RPO}}(\theta)=\frac{1}{n}\sum_{i:x_{i}=x}\delta_{i}(\theta)\cdot\nabla_{z_{\theta}(y|x)}\log\pi_{\theta}(y_{i}|x).$ (24)
@@
$\footnotesize\nabla_{z_{\theta}(y|x)}\mathcal{L}_{\mathrm{RPO}}(\theta)=\frac{1}{n}\sum_{i:x_{i}=x}\delta_{i}(\theta)\left(\mathbf{1}_{y=y_{i}}-\pi_{\theta}(y|x)\right).$ (26)
@@ IV-A Dataset and Base Models
[…]
We adopt the Flan-T5 model family [28], a series of encoder-decoder transformers fine-tuned for instruction following. To examine the effect of model scale on RPO performance, we experiment with three variants: Flan-T5 Small (77M parameters), Flan-T5 Large (783M parameters), and Flan-T5 XL (2.85B parameters). For clarity, we refer to these models as Small, Large, and XL, respectively, throughout the remainder of the paper.
All models are initialized from instruction-tuned checkpoints following the FLAN recipe [29], serving as the reference policy $\pi_{\text{ref}}$. This consistent initialization allows for direct and fair comparison of RPO across model sizes under the same training framework.
@@ IV-B Evaluation
To assess model performance, we construct a validation set of 1,000 examples from the UltraFeedback dataset [27]. These examples are selected based on their high reward scores (equal to 10.0), which indicates strong user preference and high-quality completions. This curated subset serves as the basis for two complementary evaluation methods.
Automatic Metric-Based Validation. The first evaluation strategy is conducted during training to monitor model progression. We periodically evaluate the model outputs using standard automatic metrics, including BLEU, ROUGE-L, METEOR, BERTScore, SacreBLEU, and Toxicity. These metrics—commonly used in text generation—provide surface-level assessments of similarity and relevance to reference completions. While useful for tracking learning dynamics and verifying general quality trends, such metrics are limited in capturing deeper alignment, preference satisfaction, or nuanced helpfulness.
LLM-Based Preference Evaluation. To address the limitations of automatic metrics in the reinforcement learning with human feedback (RLHF) setting, we adopt a second evaluation method based on pairwise comparisons judged by large language models (LLMs). This approach, commonly referred to as LLM-as-a-judge [30], provides a more human-aligned assessment of response quality. After training, we compare model completions using state-of-the-art LLMs—Mistral v3 [31] and LLaMA v3 [32]—as impartial judges.
Each evaluation instance presents the judge with an instruction and two responses (A and B), and the model is asked to decide which response better fulfills the instruction in terms of helpfulness and relevance. If neither response is clearly better—or if both are unhelpful or irrelevant—the model is instructed to choose "Neutral".
The prompt presented to the LLM judge is as follows:
You are an impartial assistant that judges responses based on helpfulness and relevance.
You are provided with an instruction and two responses, A and B.
Decide which response better fulfills the instruction in terms of helpfulness and relevance.
If both responses are equally good or equally poor, or if neither is clearly better, answer "Neutral".
Answer strictly with "A", "B", or "Neutral". Do NOT provide explanations or any additional text.
Instruction:
{instruction}
Response A:
{response_a}
Response B:
{response_b}
Your answer (A, B, or Neutral):
This evaluation setup allows for more robust and preference-sensitive comparisons, better reflecting human feedback and alignment goals in RLHF settings. By explicitly permitting neutral responses, we reduce bias from forced binary choices and better capture nuanced or inconclusive cases.
@@ IV-C Training Setup and Baseline Comparison
We train all models using the AdamW [33] optimizer with a learning rate of $1\times 10^{-4}$. The learning rate scheduler follows a linear warmup and decay schedule, with 150 warmup steps and total training steps computed as the product of steps per epoch and the number of epochs:
[…]
The scheduler is configured using the get_scheduler function from the Hugging Face Transformers library with linear scheduling.
For evaluating the hyperparameters of our proposed method RPO, we use KTO and DRO as baselines. The temperature parameter $\tau$ is set to 1.0 in all experiments to minimize its impact and focus on intrinsic method differences. Setting $\tau=1.0$ ensures a fair comparison by controlling the sharpness of the policy distribution uniformly.
Figure 1 compares the average training durations in hours across three FLAN-T5 model sizes (Small, Large, XL) for KTO, DRO, and RPO.
@@
As illustrated, RPO converges faster than both KTO and DRO for all model scales, highlighting its efficiency and suitability for scalable single-trajectory policy optimization.
@@ IV-D Automatic Metric-Based Validation
During training, we perform automatic validation at various checkpoints using standard natural language generation metrics to evaluate the quality of the generated responses. This validation protocol provides quantitative feedback on the model’s performance in terms of semantic similarity, fluency, and toxicity. We compare three training methods: KTO, DRO, and our proposed method RPO across three model sizes: Small, Large, and XL.
The metrics used include BERTScore F1, ROUGE-L, BLEU, SacreBLEU, METEOR, and average toxicity. Arrows indicate the desired direction for each metric: $\uparrow$ means higher values are better, while $\downarrow$ means lower values are better.
@@
TABLE II: Automatic metric-based validation results for KTO, DRO, and our proposed RPO method across different model sizes. Metrics with $\uparrow$ indicate that higher values are better, while $\downarrow$ indicates lower values are preferable. RPO consistently achieves better scores in most metrics, highlighting its effectiveness.
Table II summarizes the validation metrics across different training methods and model scales. We observe that RPO generally outperforms both KTO and DRO in semantic similarity and fluency metrics while maintaining lower toxicity levels. This suggests that RPO enhances the preference alignment of generated responses more effectively than baseline approaches.
However, these standard metrics have inherent limitations in the context of RLHF tasks. Metrics like BLEU and ROUGE focus on n-gram overlap and may fail to capture deeper semantic relevance or instruction adherence. Toxicity scores, while useful, do not fully represent nuanced safety concerns. Consequently, relying solely on these metrics can be misleading when optimizing for user preferences.
To address these limitations, we complement this validation with human-like preference evaluations conducted by large language model (LLM) judges, as described in the following subsection.
@@ IV-E LLM-Based Preference Evaluation
To better evaluate alignment with human preferences, we use two strong LLM-based judges—Mistral and LLaMA—to assess outputs from our proposed RPO method compared to three baselines: SFT, KTO, and DRO. We report the percentage of generations where RPO is preferred, along with standard deviation, across three model scales: Small, Large, and XL.
Model Size SFT KTO DRO
Small 82.78 $\pm$ 1.93 82.60 $\pm$ 2.39 83.97 $\pm$ 2.34
Large 84.20 $\pm$ 1.64 86.53 $\pm$ 1.77 84.86 $\pm$ 2.05
XL 83.01 $\pm$ 2.39 93.62 $\pm$ 5.88 83.19 $\pm$ 0.88
(a) Judgment by Mistral
Model Size SFT KTO DRO
Small 63.66 $\pm$ 1.62 65.76 $\pm$ 0.59 65.26 $\pm$ 0.41
Large 73.02 $\pm$ 0.52 72.30 $\pm$ 1.72 74.81 $\pm$ 0.63
XL 68.43 $\pm$ 1.17 61.18 $\pm$ 9.33 67.15 $\pm$ 3.60
(b) Judgment by LLaMA
TABLE III: Winrate (%) of RPO over baseline methods (SFT, KTO, DRO) using LLM judges. Higher values indicate stronger preference for RPO completions.
Analysis. Table III shows that RPO achieves consistently higher winrates than all baselines across sizes and judges. Under the more discriminative Mistral judge, RPO achieves especially strong margins, often exceeding 80% preference.
Why RPO Wins. KTO lacks explicit value modeling and learns from ranking alone, which limits its generalization. DRO introduces value estimation via a dual-network setup (value + policy), but this increases training complexity and potential instability. In contrast, RPO directly optimizes policy through preference ranking, avoiding the need to model or estimate value explicitly. This simplicity leads to more stable training and stronger performance across scales without additional modeling overhead.
@@ IV-F Impact of Regularisation Parameter $\tau$
The temperature parameter $\tau$ in Algorithm 1 plays a key role in shaping the softness of the implicit value estimation $\widehat{V}(x)$, acting as a regularisation factor between reward signals and policy logits. Lower values of $\tau$ emphasize sharper preferences (more deterministic), while higher values smooth the learned policy’s alignment with reference preferences.
We ablate the impact of $\tau$ by evaluating RPO against SFT for $\tau\in\{0.1,1.0,5.0\}$ using LLM judges (Mistral and LLaMA). The winrate of RPO over SFT is reported in Table IV. Results at $\tau=1.0$ correspond to those in Section IV-E.
Model Size $\boldsymbol{\tau=0.1}$ $\boldsymbol{\tau=1.0}$ $\boldsymbol{\tau=5.0}$
Small 84.45 $\pm$ 1.14 82.78 $\pm$ 1.93 83.03 $\pm$ 1.55
Large 59.50 $\pm$ 19.62 84.20 $\pm$ 1.64 83.72 $\pm$ 1.05
XL 90.24 $\pm$ 6.66 83.01 $\pm$ 2.39 82.81 $\pm$ 3.76
(a) Judgment by Mistral
Model Size $\boldsymbol{\tau=0.1}$ $\boldsymbol{\tau=1.0}$ $\boldsymbol{\tau=5.0}$
Small 78.73 $\pm$ 1.07 63.66 $\pm$ 1.62 66.52 $\pm$ 1.55
Large 56.40 $\pm$ 12.21 73.02 $\pm$ 0.52 74.86 $\pm$ 1.63
XL 82.07 $\pm$ 5.30 68.43 $\pm$ 1.17 66.28 $\pm$ 2.48
(b) Judgment by LLaMA
TABLE IV: Winrate (%) of RPO over SFT across different values of the regularization parameter $\tau$, measured by pairwise preference judgment.
Interpretation. Table IV highlights that RPO is generally robust to $\tau$ variation, particularly at $\tau=1.0$ and $\tau=5.0$. While $\tau=0.1$ performs well at Small and XL scales, it suffers a drop at the Large scale—especially under the Mistral judge—likely due to instability from overly sharp preference weighting.
Moderate values ($\tau=1.0$) yield consistent and strong results, balancing stability and preference sensitivity. Larger $\tau$ values retain strong performance, showing that RPO does not require precise tuning of the temperature. This is a notable advantage compared to methods that tightly couple policy learning with value estimation or reward calibration.
This supports our core insight: RPO avoids direct value modeling, estimating $\widehat{V}(x)$ implicitly via a reference-weighted partition function. It focuses training entirely on policy logits, sidestepping the instability often observed in methods like DRO which rely on dual-network learning (policy + value), and simplifying optimization compared to actor-critic or reward-model-based alternatives.
@@ IV-G Ablation Study: Importance of Empirical Normalization
To assess the role of empirical normalization in Reward Policy Optimization (RPO), we introduce an ablated variant, RPO-noNorm, where the baseline estimate $\widehat{V}(x)$ is removed and the raw reward $r(x,y)$ is used directly. This change eliminates advantage-based learning and removes the normalization effect that stabilizes optimization. We compare both variants across temperatures $\tau\in\{0.1,1.0,5.0\}$ and model sizes (Small, Large, XL), using both automatic and human evaluations.
Automatic Evaluation Metrics. Table V presents results across six metrics. We observe:
  • RPO consistently outperforms RPO-noNorm across all metrics, sizes, and $\tau$ values.
  • The performance of RPO is stable with respect to temperature, showing its robustness.
  • RPO-noNorm fails to benefit from temperature scaling and shows consistently lower scores.
  • Toxicity increases substantially without normalization, suggesting unstable reward learning.
@@
ROUGE-L ↑ 0.2861 0.3143 0.2951 0.2885 0.3162 0.2969 0.2872 0.3150 0.2958 0.2063 0.2028 0.2051 0.2089 0.2043 0.2067 0.2115 0.2078 0.2101
BLEU ↑ 0.0354 0.0460 0.0503 0.0360 0.0464 0.0511 0.0348 0.0456 0.0498 0.0130 0.0127 0.0139 0.0135 0.0129 0.0143 0.0131 0.0132 0.0137
SacreBLEU ↑ 3.5421 4.7650 5.0228 3.6013 4.7903 5.1063 3.5470 4.7582 5.0157 1.2360 1.2203 1.2632 1.2743 1.2312 1.2914 1.2634 1.2398 1.2665
METEOR ↑ 0.2401 0.2619 0.2551 0.2418 0.2632 0.2567 0.2403 0.2626 0.2559 0.1749 0.1694 0.1726 0.1762 0.1708 0.1736 0.1780 0.1719 0.1741
@@
TABLE V: Automatic metric comparison of RPO and RPO-noNorm across temperatures $\tau\in\{0.1,1.0,5.0\}$ and model sizes. S, L, and XL refer to Small, Large, and Extra-Large model variants, respectively. RPO consistently achieves stronger performance and lower toxicity across all configurations. Temperature $\tau$ has minimal effect compared to normalization.
These results demonstrate that empirical normalization via $\widehat{V}(x)$ is essential for robust and high-quality optimization. Without it, models suffer large drops in fluency, adequacy, and safety—effects that cannot be corrected by tuning $\tau$ alone.
Human Evaluation via LLM Judges. Table VI reports pairwise win-rates of RPO over RPO-noNorm as judged by Mistral and LLaMA. RPO is preferred in over 90% of comparisons across all $\tau$ and model sizes. This reinforces that empirical normalization not only improves automatic metrics but is also recognized as producing superior outputs by strong language models.
Model Size $\boldsymbol{\tau=0.1}$ $\boldsymbol{\tau=1.0}$ $\boldsymbol{\tau=5.0}$
Small 96.22 $\pm$ 0.91 95.04 $\pm$ 0.84 95.77 $\pm$ 1.02
Large 94.85 $\pm$ 0.77 96.13 $\pm$ 0.66 95.90 $\pm$ 0.58
XL 97.34 $\pm$ 0.52 95.81 $\pm$ 0.63 96.12 $\pm$ 0.49
(a) Judgment by Mistral
Model Size $\boldsymbol{\tau=0.1}$ $\boldsymbol{\tau=1.0}$ $\boldsymbol{\tau=5.0}$
Small 92.75 $\pm$ 1.13 91.34 $\pm$ 1.25 91.80 $\pm$ 0.97
Large 90.64 $\pm$ 1.31 91.76 $\pm$ 0.89 92.08 $\pm$ 0.74
XL 94.21 $\pm$ 1.03 92.99 $\pm$ 0.85 93.17 $\pm$ 0.77
(b) Judgment by LLaMA
TABLE VI: LLM preference win-rate (%) of RPO over RPO-noNorm across temperatures and model sizes. RPO is consistently preferred, confirming the necessity of normalization.
In conclusion, empirical normalization is a decisive factor in the effectiveness of RPO. Removing $\widehat{V}(x)$ leads to clear and consistent degradation in both automatic metrics and human preference. While temperature $\tau$ modulates exploration, it cannot substitute the stabilizing effect of normalization, making it a critical design choice in reward optimization.

### v3

@@ III Method
$\footnotesize\pi^{*}_{\theta}(y|x)=\frac{\pi_{\mathrm{ref}}(y|x)\exp\left(\frac{1}{\tau}r(x,y)\right)}{Z^{*}(x)},$ (10)
where the partition function $Z^{*}(x)$ is
$\footnotesize Z^{*}(x)=\sum_{y^{\prime}}\pi_{\mathrm{ref}}(y^{\prime}|x)\exp\left(\frac{1}{\tau}r(x,y^{\prime})\right).$ (11)
Current methods such as DRO learn both a policy $\pi_{\theta}$ and a value function $V_{\phi}$ parameterized by $\phi$, where the soft value function relates to the partition function as $V^{*}(x)=\tau\log Z^{*}(x)$. However, this joint learning introduces several difficulties: the value function $V_{\phi}$ can be unstable to approximate, the parameters $(\theta,\phi)$ are tightly coupled, and optimization becomes more complex.
@@
$\footnotesize r(x,y)-V^{*}(x)=\tau\log\frac{\pi^{*}_{\theta}(y|x)}{\pi_{\mathrm{ref}}(y|x)},$ (12)
@@
$\footnotesize\hat{\mathcal{L}}_{\mathrm{RPO}}(\theta)=\frac{1}{2N}\sum_{i=1}^{N}\left(\log\frac{\pi_{\theta}(y_{i}|x_{i})}{\pi_{\mathrm{ref}}(y_{i}|x_{i})}-\frac{1}{\tau}\left(r_{i}-\widehat{V}(x_{i})\right)\right)^{2}.$ (16)
@@
$\footnotesize\nabla_{z_{\theta}(y|x)}\hat{\mathcal{L}}_{\mathrm{RPO}}(\theta)=\frac{1}{N}\sum_{i:x_{i}=x}\delta_{i}(\theta)\cdot\nabla_{z_{\theta}(y|x)}\log\pi_{\theta}(y_{i}|x).$ (19)
@@
$\footnotesize\nabla_{z_{\theta}(y|x)}\hat{\mathcal{L}}_{\mathrm{RPO}}(\theta)=\frac{1}{N}\sum_{i:x_{i}=x}\delta_{i}(\theta)\left(\mathbf{1}_{y=y_{i}}-\pi_{\theta}(y|x)\right).$ (21)
@@ IV-A Dataset and Base Models
We follow the experimental setup of Richemond et al. [richemond2024offline] and use the UltraFeedback dataset [cui2023ultrafeedback], which consists of offline triplets $(\text{prompt},\text{completion},\text{reward})$ with multiple completions per prompt. Rewards are standardized to zero mean and unit variance for training stability.
We use instruction-tuned Flan-T5 models [roberts2023scaling] at three scales: Small (77M), Large (783M), and XL (2.85B). All models are initialized from FLAN instruction-tuned checkpoints [chung2024scaling], which serve as the reference policy $\pi_{\text{ref}}$, enabling fair comparison across model sizes.
@@ IV-B Evaluation
We evaluate models on a validation set of 1,000 high-quality examples (reward =10.0=10.0) from the UltraFeedback dataset [cui2023ultrafeedback]. This subset is used for two complementary evaluation protocols.
Automatic Metrics. During training, we monitor performance using standard generation metrics: BLEU, ROUGE-L, METEOR, BERTScore, SacreBLEU, and Toxicity. These metrics capture surface-level similarity and general quality trends but are limited in assessing alignment and preference satisfaction.
LLM-Based Preference Evaluation. To obtain more human-aligned judgments, we perform post-training pairwise evaluations using the LLM-as-a-judge paradigm [gu2024survey]. Model outputs are compared using Mistral v3 [jiang2023mistral7b] and LLaMA v3 [touvron2023llamaopenefficientfoundation] as impartial judges. Given an instruction and two responses, the judge selects the better one based on helpfulness and relevance, or Neutral if neither is clearly preferred.
Allowing neutral judgments mitigates bias from forced binary decisions and enables more reliable preference-sensitive evaluation in RLHF settings.
@@ IV-C Training Setup and Baseline Comparison
All models are trained with the AdamW optimizer [loshchilov2019decoupled] using a learning rate of $1\times 10^{-4}$. We apply a linear warmup and decay schedule with 150 warmup steps, where the total number of training steps is given by
$\texttt{num\_training\_steps}=\texttt{steps\_per\_epoch}\times\texttt{epochs}.$
We compare RPO against KTO and DRO baselines. To isolate methodological differences, the temperature parameter is fixed to $\tau=1.0$ across all experiments.
@@
Figure 1 reports average training time across FLAN-T5 Small, Large, and XL models. RPO consistently converges faster than KTO and DRO at all scales, demonstrating superior training efficiency and scalability.
@@ IV-D Automatic Metric-Based Validation
We evaluate model quality during training using standard automatic generation metrics across three methods (KTO, DRO, and RPO) and three model sizes (Small, Large, XL). The metrics include BERTScoreF1, ROUGE-L, BLEU, SacreBLEU, METEOR, and average toxicity, where $\uparrow$ indicates higher-is-better and $\downarrow$ lower-is-better.
@@
TABLE II: Automatic metric-based validation results for KTO, DRO, and RPO across different model sizes.
Table II shows that RPO consistently outperforms KTO and DRO across semantic similarity and fluency metrics while achieving lower toxicity at all model scales. These results indicate stronger preference alignment, though automatic metrics alone are insufficient to fully capture human judgments. We therefore complement this analysis with LLM-based preference evaluation in the following subsection.
@@ IV-E LLM-Based Preference Evaluation
We evaluate preference alignment using two LLM judges, Mistral and LLaMA, comparing RPO against SFT, KTO, and DRO across Small, Large, and XL models. We report the win rate (%) of RPO over each baseline, with standard deviation.
Model Size Mistral Judge LLaMA Judge
SFT KTO DRO SFT KTO DRO
Small 82.78 $\pm$ 1.93 82.60 $\pm$ 2.39 83.97 $\pm$ 2.34 63.66 $\pm$ 1.62 65.76 $\pm$ 0.59 65.26 $\pm$ 0.41
Large 84.20 $\pm$ 1.64 86.53 $\pm$ 1.77 84.86 $\pm$ 2.05 73.02 $\pm$ 0.52 72.30 $\pm$ 1.72 74.81 $\pm$ 0.63
XL 83.01 $\pm$ 2.39 93.62 $\pm$ 5.88 83.19 $\pm$ 0.88 68.43 $\pm$ 1.17 61.18 $\pm$ 9.33 67.15 $\pm$ 3.60
TABLE III: Win rate (%) of RPO over SFT, KTO, and DRO as judged by Mistral and LLaMA. Higher is better.
Table III shows that RPO consistently outperforms all baselines across model sizes and judges, with particularly strong margins under the more discriminative Mistral judge. These results indicate that RPO produces responses better aligned with human preferences than competing methods.
RPO’s advantage stems from directly optimizing the policy via preference ranking, avoiding explicit value modeling required by KTO and DRO. This simpler optimization leads to more stable training and improved preference alignment across scales.
@@ IV-F Impact of Regularisation Parameter $\tau$
We study the effect of the temperature parameter $\tau$ in RPO by comparing its win rate over SFT for $\tau\in\{0.1,1.0,5.0\}$ using Mistral and LLaMA judges. Results for $\tau=1.0$ match those reported in Section IV-E.
Model Size Mistral Judge LLaMA Judge
$\tau=0.1$ $\tau=1.0$ $\tau=5.0$ $\tau=0.1$ $\tau=1.0$ $\tau=5.0$
Small 84.45 $\pm$ 1.14 82.78 $\pm$ 1.93 83.03 $\pm$ 1.55 78.73 $\pm$ 1.07 63.66 $\pm$ 1.62 66.52 $\pm$ 1.55
Large 59.50 $\pm$ 19.62 84.20 $\pm$ 1.64 83.72 $\pm$ 1.05 56.40 $\pm$ 12.21 73.02 $\pm$ 0.52 74.86 $\pm$ 1.63
XL 90.24 $\pm$ 6.66 83.01 $\pm$ 2.39 82.81 $\pm$ 3.76 82.07 $\pm$ 5.30 68.43 $\pm$ 1.17 66.28 $\pm$ 2.48
TABLE IV: Win rate (%) of RPO over SFT for different values of $\tau$, evaluated by Mistral and LLaMA.
Table IV shows that RPO is robust to the choice of $\tau$, with strong and stable performance at $\tau=1.0$ and $\tau=5.0$ across model sizes and judges. Very low temperature ($\tau=0.1$) can introduce instability at larger scales, particularly under Mistral.
Overall, these results indicate that RPO does not require precise temperature tuning, supporting its practicality and stability compared to methods that rely on explicit value modeling or tightly coupled reward calibration.
@@ IV-G Ablation Study: Importance of Empirical Normalization
We study the role of empirical normalization in RPO by comparing the full method against an ablated variant, RPO-noNorm, which removes the baseline $\widehat{V}(x)$ and directly optimizes raw rewards. We evaluate both methods across $\tau\in\{0.1,1.0,5.0\}$ and model sizes (Small, Large, XL).
Automatic Metrics. Table V shows that RPO consistently outperforms RPO-noNorm across all metrics, temperatures, and model scales. RPO remains stable under temperature variation and benefits from scaling, whereas RPO-noNorm exhibits degraded fluency, weaker semantic alignment, higher toxicity, and minimal gains from larger models.
@@
ROUGE-L 0.2861 0.3143 0.2951 0.2885 0.3162 0.2969 0.2872 0.3150 0.2958 0.2063 0.2028 0.2051 0.2089 0.2043 0.2067 0.2115 0.2078 0.2101
BLEU 0.0354 0.0460 0.0503 0.0360 0.0464 0.0511 0.0348 0.0456 0.0498 0.0130 0.0127 0.0139 0.0135 0.0129 0.0143 0.0131 0.0132 0.0137
SacreBLEU 3.5421 4.7650 5.0228 3.6013 4.7903 5.1063 3.5470 4.7582 5.0157 1.2360 1.2203 1.2632 1.2743 1.2312 1.2914 1.2634 1.2398 1.2665
METEOR 0.2401 0.2619 0.2551 0.2418 0.2632 0.2567 0.2403 0.2626 0.2559 0.1749 0.1694 0.1726 0.1762 0.1708 0.1736 0.1780 0.1719 0.1741
@@
TABLE V: Automatic metric comparison of RPO and RPO-noNorm across temperatures and model sizes.
LLM-Based Preference Evaluation. As shown in Table VI, RPO is preferred over RPO-noNorm in over 90% of comparisons across all temperatures and model sizes, under both Mistral and LLaMA judges.
Model Size Mistral LLaMA
$\tau=0.1$ $\tau=1.0$ $\tau=5.0$ $\tau=0.1$ $\tau=1.0$ $\tau=5.0$
Small 96.22 $\pm$ 0.91 95.04 $\pm$ 0.84 95.77 $\pm$ 1.02 92.75 $\pm$ 1.13 91.34 $\pm$ 1.25 91.80 $\pm$ 0.97
Large 94.85 $\pm$ 0.77 96.13 $\pm$ 0.66 95.90 $\pm$ 0.58 90.64 $\pm$ 1.31 91.76 $\pm$ 0.89 92.08 $\pm$ 0.74
XL 97.34 $\pm$ 0.52 95.81 $\pm$ 0.63 96.12 $\pm$ 0.49 94.21 $\pm$ 1.03 92.99 $\pm$ 0.85 93.17 $\pm$ 0.77
TABLE VI: LLM preference win rate (%) of RPO over RPO-noNorm.
These results demonstrate that empirical normalization via $\widehat{V}(x)$ is critical for stable and effective optimization. Temperature tuning alone cannot compensate for its removal, confirming normalization as a key component of RPO.